A 20 acre field yields 610 bushels of corn. How much would a 25 acre field yield under similar conditions?
Answer by tuxrocks(3)

610/4=152.50
152.50*5=762.50
A 25 acre field would yeild 762.50 bushels of corn.
Imagine there are 2 fields. Field 1 has 4 subfields. Field 2 has 5 subfields. Field 1 has 610 bushels. Field 2 has how many bushels? 610/4=152.50 bushels per subfield. 152.50(bushels per subfield) times 5(number of field 2 subfields)=762.50